The pH 10 buffer used involves mixing ammonia and ammonium chloride. What is the molar ratio of ammonia to ammonium chloride needed to produce a buffer with a pH of 10?
1
Expert's answer
2019-03-18T02:29:07-0400

To find pH of the ammonia buffer we use the formula

.pH=14-p+lg

.lg = pH-14+p=10-14+4.75=0.75

=1=5.6